How many liters of air will burn 30 liters of methane?
1
Expert's answer
2017-11-01T14:10:07-0400
〖CH〗_4+〖2O〗_2→〖CO〗_2+〖2H〗_2 O
ν=V/V_m ⇒ν~V
V(O_2 )=2V(〖CH〗_4 )=30∙2=60 l.
Air contains 21% O_2 by volume.
V(air)=V(O_2 )/0.21=285.7≈286 l.
Answer
286 l.